# Unsigned Orthogonal Distance Fields: An Accurate Neural Implicit Representation for Diverse 3D Shapes

### v1

@@ 1 Introduction
In this paper, we propose the unsigned orthogonal distance fields (UODFs) based NIR. Compared with conventional SDF and UDF, the most distinguishable feature of UODFs is that the minimal distances to the shape surface are defined along three orthogonal directions in 3D space. This means every point in UODFs can directly access its closest surface points along three orthogonal directions, when they exist. Fig. 1 (a) depicts how SDF (or UDF) and UODFs based NIR work in surface points reconstruction. All predicted distance values denoted by $\widehat{...D\!F(s)}$ at grid corners are affected by the fitting error of neural network. However, in contrast to SDF or UDF, the estimation of GEP for UODFs does not introduce interpolation errors. Instead, the fitting error can be mitigated through averaging multiple predictions for each GEP. Therefore, UODFs based NIR naturally leverages accurate reconstruction for diverse 3D shapes, three of which are shown in Fig. 1 (b).
@@ 2 Related Works
The focus in this paper lies in the neural distance field representation of 3D shapes. This research topic could be broadly grouped into three categories for review: SDF based NIR, UDF based NIR, and other types of neural distance field representation.
@@ 3.2 Characteristics of UODFs
Characteristic 3. $U\!O\!D\!F_{X}$ may display discontinuity between adjacent parallel rays, which distinguishes it from the SDF or UDF that is continuous everywhere. Fig. 3 illustrates a slice of $U\!O\!D\!F_{U\!D}$ from the 3D model named ‘Dragon’. It is evident that the $U\!O\!D\!F_{U\!D}$ values of two neighborly points on adjacent rays can differ greatly. This unique characteristic mirrors the behavior observed in laser sensors and facilitates the direct and accurate prediction of fine details.
@@ 3.3 Network Architecture and Processing Pipeline
Fig. 4 depicts the proposed network architecture and processing pipeline. Parallel rays are defined on each orthogonal plane. For a spatial point, its 2D coordinates on one orthogonal plane are employed to extract a $42$-dimensional ray feature via position encoding (‘P. E.’ module). This ray feature, combined with the original 3D coordinates, forms the $45$-dimensional input for a multi-layer perceptron (MLP) network to regress the UODF in this orthogonal direction. Meanwhile, to avoid regressing the UODF of non-intersected rays which are not defined in Eq. 2, the ray feature is also utilized to regress a 2D mask on this orthogonal plane (see the area in each silhouette). Subsequently, the surface points in this orthogonal direction can be estimated from rays that intersect the shape, which will be illustrated in Sec. 3.5.
The networks to regress the UODF and the mask in each orthogonal direction are a $10$***$256$ MLP and $3$***$256$ MLP, respectively. The supervised mask values are set to be ‘0’ and ‘1’, denoting the outside and inside of the silhouette respectively. The loss function of $U\!O\!D\!F_{X}$ is composed of three component loss terms:
@@ 3.6 Mesh Extraction
Despite the SPSR method, without the fine-tuning of the default hyper-parameters, generates superior mesh results in our experiments, we explore another way for simultaneous GEP reconstruction and mesh extraction, based on the MeshUDF [14] algorithm. To summarise, we compute the normal vectors and UDF values at grid corners from our predicted UODFs and feed them into MeshUDF as an input. Detailed process and experimental results are available in the supplementary material.
@@ 4.1 Dataset and Metrics
[…]
@@ 4.3 Reconstruction of Non-watertight Shapes
CD-GEP (*$10^{5}$$\bm{\downarrow}$ 128 13.0 4.95 0.227
CD-Mesh (*$10^{5}$$\bm{\downarrow}$ 128 14.4 6.1 1.93
NC-Mesh $\bm{\downarrow}$ 90.8 88.0 92.0 99.6
@@ A.4 Parameter Amount
Table A.1 presents a comprehensive comparison of the total number of parameters utilized in various methods. In our method, we employ a configuration of three $10$***$256$ and $3$***$256$ MLPs to separately regress UODF in each orthogonal direction. Despite this seemingly large network setup, our parameter amount (1.84M) is notably smaller than those in three UDF regression methods and the state-of-the-art SDF regression method NGLOD5.
@@ Appendix C Additional Details of UODFs Fitting and Reconstruction
  • The fitting and reconstruction results at $257^{3}$ grids are displayed in Fig. C.2. The first row depicts the predicted masks on three orthogonal planes, accompanied by the number of rays (NoR) in each mask.
@@
These instances exemplify the effectiveness of our UODFs fitting and reconstruction process. We emphasize again that the experimental results in our manuscript and video are free of fine-tuning, although this can further improve the visual effect and accuracy metrics.
@@
Figure C.3: One slice of the $U\!O\!D\!F_{L\!R}$ and $U\!O\!D\!F_{F\!B}$ of the ‘Dragon’ model. A subtle difference between ground truth and our prediction can be seen in the areas drawn by black dotted circles (with zoomed-in views).
@@ Appendix D Ablation of Point Merging Threshold
Appendix E Alternative Reconstruction Approach Employing MeshUDF
In this section, we introduce an alternative approach for reconstruction, employing the MeshUDF method [14]. MeshUDF, extending MC to extract mesh from UDF, requires the gradients and values of UDF at grid corners as input. In alignment with MeshUDF’s requirements, our alternative approach deviates from the presented practice of estimating GEP from rays followed by meshing, illustrated in the main manuscript. Instead, we explore estimating the gradients and values of UDF directly from the predicted UODFs at grid corners.
E.1 Gradient Computation
The process of computing the gradients of UODFs is shown in Fig. D.4. Consider a local triangular patch near a grid corner, denoted as $O$. The gradient directions of UDF and each UODF are the same, pointing perpendicular to the triangular patch. The upper left part in the figure details the computation of the three gradient components for a single UODF, denoted as $U\!O\!D\!F_{X}$ along the $x$ direction. Specifically, $U\!O\!D\!F_{X}(O)$ indicates the distance from the corner $O$ to the $x$-axis intersection with the triangular patch. The partial derivatives are calculated by determining the $U\!O\!D\!F_{X}$ values at slightly shifted positions along the $x$, $y$, and $z$ axes. It can be seen that the computed gradient $grad_{X}$ aligns with the normal of the triangular patch, as depicted in the lower left of the figure. Similarly, the gradients for the other two UODFs, $grad_{Y}$ and $grad_{Z}$ theoretically coincide with $grad_{X}$ and represent the triangle’s normal. Considering practical errors, the final gradient is obtained by fusing these three calculated gradients, as illustrated in the figure’s right side.
E.2 UDF Computation
Regarding the UDF computation for this simple local shape, the UDF value corresponds to the distance from the corner $O$ to the triangular patch. Thus, $U\!O\!D\!F_{X}(O)$ can be projected onto the unit vector of the fused gradient to derive this distance. Likewise, we can obtain three distance values from the UODFs values along the orthogonal directions. Among these three calculated distances, we select the smallest one as the final UDF result, as the minimum distance most accurately indicates the proximity of the triangular patch to the corner.
Figure E.5: Reconstruction results obtained by applying the described approach to estimate gradients and UDF values at grid corners, followed by reconstruction employing the MeshUDF algorithm.
E.3 Results Presentation
Fig. E.5 presents the results achieved using our specialized method for calculating gradients and UDF values at grid corners, subsequently employing the MeshUDF algorithm for reconstruction. While most reconstructed meshes are visually satisfactory, they occasionally exhibit issues such as reduced accuracy and robustness. Therefore, we advocate for the use of screened Poisson surface reconstruction (SPSR) [18], as introduced in the main manuscript, for post-processing reconstruction. This recommendation is grounded in SPSR’s ability to effectively address these specific deficiencies, enhancing both the accuracy and completeness of the meshes.
Appendix F Additional Reconstruction Results
Table F.4: Detailed reconstruction metrics for each of ten garments from the MGN dataset. The two CD metrics $\bm{\downarrow}$ are multiplied by a factor of $10^{5}$.
@@ Appendix F Additional Reconstruction Results
First, we elaborate on the three reconstruction metrics for the ten non-watertight garments from the MGN dataset in Table F.4. These results show that our UODFs based NIR method significantly outperforms three other UDF regression methods. We have already showcased the ‘TshirtNoCoat-1’ garment in Fig. 9 in the main manuscript and the results for the remaining nine garments are illustrated in Fig.F.6.
Second, for the watertight shapes, additional results on the Thingi32 dataset are displayed in Fig. F.7, demonstrating consistency with the findings shown in Fig. 6 and Table 1 in the main manuscript.
Third, we present the outcomes for the additional four complex shapes in Fig. F.8. The zoomed-in views further underscore the high-quality shape reconstruction achievable with our proposed UODFs.
Finally, more reconstruction experiments at various grid resolutions from $32^{3}$ to $256^{3}$ are conducted. Fig. F.9 shows the results of four distinct shapes: two watertight and two non-watertight. To highlight the reconstruction accuracy of our UODF-based NIR method, we compare it with the marching cube (MC) algorithm [23] and MeshUDF [14], that utilize ground truth SDF and UDF as input, respectively. Therefore, the MC and MeshUDF results do not suffer from fitting errors of neural network, but are only affected by interpolation errors. For these two watertight models, our method achieves performance comparable to the MC method across all grid resolutions. In the case of non-watertight models, our method significantly outperforms MeshUDF using the ground truth UDF as input, at all grid resolutions. MeshUDF has some spurious meshes at the edge of shapes, especially under lower resolutions. In addition, our meshes generated by SPSR [18] are smoother than those extracted by MC and MeshUDF.
Figure F.6: Reconstruction results for other nine shapes in the MGN10 dataset.
Figure F.7: Reconstruction results for more shapes in the Thingi32 dataset. The metrics from left to right below each shape are CD-GEP(*$10^{5}$$\bm{\downarrow}$, CD-Mesh(*$10^{5}$$\bm{\downarrow}$, and NC-Mesh $\bm{\uparrow}$, respectively.
Figure F.8: Reconstruction results for additional four complex shapes.
Figure F.9: Additional reconstruction results at various resolutions of grids. The metrics from left to right below each shape are CD-GEP(*$10^{5}$$\bm{\downarrow}$, CD-Mesh(*$10^{5}$$\bm{\downarrow}$, and NC-Mesh $\bm{\uparrow}$, respectively.

### v2

@@ 1 Introduction
In this paper, we propose the unsigned orthogonal distance fields (UODFs) based NIR. Compared with conventional SDF and UDF, the most distinguishable feature of UODFs is that the minimal distances to the shape surface are defined along three orthogonal directions in 3D space. This means every point in UODFs can directly access its closest surface points along three orthogonal directions, when they exist. Fig. 1 (a) depicts how SDF (or UDF) and UODFs based NIR work in surface points reconstruction. All predicted distance values denoted by $\footnotesize\widehat{...D\!F(s)}$ at grid corners are affected by the fitting error of neural network. However, in contrast to SDF or UDF, the estimation of GEP for UODFs does not introduce interpolation errors. Instead, the fitting error can be mitigated through averaging multiple predictions for each GEP. Therefore, UODFs based NIR naturally leverages accurate reconstruction for diverse 3D shapes, three of which are shown in Fig. 1 (b).
@@ 2 Related Works
The focus in this paper lies in the neural distance field representation of 3D shapes, which could be broadly grouped into three categories for review: SDF based NIR, UDF based NIR, and other types of neural distance field.
@@ 3.2 Characteristics of UODFs
Characteristic 3. $U\!O\!D\!F_{X}$ may display discontinuity between adjacent parallel rays, which distinguishes it from the SDF or UDF that is continuous everywhere. Fig. 3 depicts a slice of $U\!O\!D\!F_{U\!D}$ from the 3D model ‘Dragon’. It is evident that the $U\!O\!D\!F_{U\!D}$ values of two neighborly points on adjacent rays can differ greatly. This unique characteristic mirrors the behavior observed in laser sensors and facilitates the direct and accurate prediction of fine details.
@@ 3.3 Network Architecture and Processing Pipeline
Fig. 4 depicts the proposed network architecture and processing pipeline. Parallel rays are defined on each orthogonal plane. For a spatial point, its 2D coordinates on one orthogonal plane are employed to extract a $42$-dimensional feature via position encoding (‘P. E.’ module). This ray feature, combined with the 3D coordinates, forms the $45$-dimensional input for a multi-layer perceptron (MLP) network to regress the UODF in this orthogonal direction. Meanwhile, to avoid regressing the UODF of non-intersected rays which are not defined in Eq. 2, the ray feature is utilized to regress a 2D mask on this orthogonal plane (see the area in each silhouette). Subsequently, the surface points in this orthogonal direction can be estimated from rays that intersect the shape, which is illustrated in Sec. 3.5.
The networks to regress the UODF and the mask in each orthogonal direction are a $10$*$256$ MLP and $3$*$256$ MLP, respectively. The supervised mask values are set to be ‘0’ and ‘1’, denoting the outside and inside of the silhouette respectively. The loss function of $U\!O\!D\!F_{X}$ is composed of three component loss terms:
@@ 3.6 Mesh Extraction
Despite the SPSR method, without the fine-tuning of the default hyper-parameters, generates superior mesh results in our experiments, we explore another way for simultaneous GEP reconstruction and mesh extraction, based on the MeshUDF [14] algorithm. These findings are preliminary and will be refined in subsequent research.
@@ 4.1 Dataset and Metrics
Metrics. The reconstruction accuracy of grid edge points (GEP) is evaluated by a $\textit{L}_{2}$ Chamfer distance (CD), abbreviated as ‘GEP-CD’. For mesh, a $\textit{L}_{2}$ CD and normal consistency (NC) of $100,000$ evenly distributed points are measured, abbreviated as ‘Mesh-CD’ and ‘Mesh-NC’ respectively. For the sake of fair comparison, all reconstructed meshes from tested NIR methods employ predicted distance values of $257^{3}$ grid corners, except the experiments with multiple grid resolutions as shown in Sec. 4.6.
@@ 4.3 Reconstruction of Non-watertight Shapes
CD-GEP (*$10^{5}$)  $\bm{\downarrow}$ 128 13.0 4.95 0.227
CD-Mesh (*$10^{5}$)  $\bm{\downarrow}$ 128 14.4 6.1 1.93
NC-Mesh $\bm{\uparrow}$ 90.8 88.0 92.0 99.6
@@ A.4 Parameter Amount
Table A.1 presents a comprehensive comparison of the total number of parameters utilized in various methods. In our method, we employ a configuration of three $10$*$256$ and $3$*$256$ MLPs to separately regress UODF in each orthogonal direction. Despite this seemingly large network setup, our parameter amount (1.84M) is notably smaller than those in three UDF regression methods and the state-of-the-art SDF regression method NGLOD5.
@@ Appendix C Additional Details of UODFs Fitting and Reconstruction
  • The fitting and reconstruction results at $257^{3}$ grid corners are displayed in Fig. C.2. The first row depicts the predicted masks on three orthogonal planes, accompanied by the number of rays (NoR) in each mask.
@@
These instances exemplify the effectiveness of our UODFs fitting and reconstruction process. We emphasize that the experimental results in our manuscript and video are free of fine-tuning, although this can further improve the visual effect and accuracy metrics.
@@
Figure C.3: One slice of the $U\!O\!D\!F_{L\!R}$ and $U\!O\!D\!F_{F\!B}$ of the ‘Dragon’ model. A subtle difference between ground truth and our prediction can be seen in the areas drawn by black dotted circles in the $U\!O\!D\!F_{F\!B}$ (with zoomed-in views).
@@ Appendix D Ablation of Point Merging Threshold
Appendix E Additional Reconstruction Results
Table E.4: Detailed reconstruction metrics for each of ten garments from the MGN dataset. The two CD metrics $\bm{\downarrow}$ are multiplied by a factor of $10^{5}$.
@@ Appendix E Additional Reconstruction Results
First, we elaborate on the three reconstruction metrics for the ten non-watertight garments from the MGN dataset in Table E.4. These results show that our UODFs based NIR method significantly outperforms three other UDF regression methods. We have already showcased the ‘TshirtNoCoat-1’ garment in Fig. 9 in the main manuscript and the results for the remaining nine garments are illustrated in Fig. E.4.
Second, for the watertight shapes, additional results on the Thingi32 dataset are displayed in Fig. E.5, demonstrating consistency with the findings shown in Fig. 6 and Table 1 in the main manuscript.
Third, we present the outcomes for the additional four complex shapes in Fig. E.6. The zoomed-in views further underscore the high-quality shape reconstruction achievable with our proposed UODFs.
Finally, more reconstruction experiments at various grid resolutions from $32^{3}$ to $256^{3}$ are conducted. Fig. E.7 shows the results of four distinct shapes: two watertight and two non-watertight. To highlight the reconstruction accuracy of our UODF-based NIR method, we compare it with the marching cube (MC) algorithm [23] and MeshUDF [14], that utilize ground truth SDF and UDF as input, respectively. Therefore, the MC and MeshUDF results do not suffer from fitting errors of neural network, but are only affected by interpolation errors. For these two watertight models, our method achieves performance comparable to the MC method across all grid resolutions. In the case of non-watertight models, our method significantly outperforms MeshUDF using the ground truth UDF as input, at all grid resolutions. MeshUDF has some spurious meshes at the edge of shapes, especially under lower resolutions. In addition, our meshes generated by SPSR [18] are smoother than those extracted by MC and MeshUDF.
Figure E.4: Reconstruction results for other nine shapes in the MGN10 dataset.
Figure E.5: Reconstruction results for more shapes in the Thingi32 dataset. The metrics from left to right below each shape are CD-GEP(*$10^{5}$$\bm{\downarrow}$, CD-Mesh(*$10^{5}$$\bm{\downarrow}$, and NC-Mesh $\bm{\uparrow}$, respectively.
Figure E.6: Reconstruction results for additional four complex shapes.
Figure E.7: Additional reconstruction results at various resolutions of grids. The metrics from left to right below each shape are CD-GEP(*$10^{5}$$\bm{\downarrow}$, CD-Mesh(*$10^{5}$$\bm{\downarrow}$, and NC-Mesh $\bm{\uparrow}$, respectively.